# STORM: A Spatio-Temporal Factor Model Based on Dual Vector Quantized Variational Autoencoders for Financial Trading

### v3

@@ Abstract.
In financial trading, factor models are widely used to price assets and capture excess returns from mispricing. Recently, we have witnessed the rise of variational autoencoder-based latent factor models, which learn latent factors self-adaptively. While these models focus on modeling overall market conditions, they often fail to effectively capture the temporal patterns of individual stocks. Additionally, representing multiple factors as single values simplifies the model but limits its ability to capture complex relationships and dependencies. As a result, the learned factors are of low quality and lack diversity, reducing their effectiveness and robustness across different trading periods. To address these issues, we propose a Spatio-Temporal factOR Model based on dual vector quantized variational autoencoders, named STORM, which extracts features of stocks from temporal and spatial perspectives, then fuses and aligns these features at the fine-grained and semantic level, and represents the factors as multi-dimensional embeddings. The discrete codebooks cluster similar factor embeddings, ensuring orthogonality and diversity, which helps distinguish between different factors and enables factor selection in financial trading. To show the performance of the proposed factor model, we apply it to two downstream experiments: portfolio management on two stock datasets and individual trading tasks on six specific stocks. The extensive experiments demonstrate STORM’s flexibility in adapting to downstream tasks and superior performance over baseline models. Code is available in PyTorch [^1] https://anonymous.4open.science/r/Storm-2823.
@@
conference: Proceedings of the 31th ACM SIGKDD Conference on Knowledge Discovery and Data Mining 2025; August 03–August 07, 2025; Toronto, Canada
@@ 1. Introduction
In financial trading, factor models are fundamental tools for asset pricing, widely used to predict asset returns. These models enhance the accuracy of asset pricing and risk management by identifying a set of key factors that explain excess returns. The well-known Fama-French three-factor model (Fama and French, 1992), for example, builds on the Capital Asset Pricing Model (Sharpe, 1964) by introducing additional factors to capture the risk premiums associated with different types of stocks in the market. As financial theories have evolved, researchers have continued to discover new factors, allowing for a more comprehensive understanding of market dynamics.
@@
Recently, we have witnessed the rise of latent factor models (Gu et al., 2021; Duan et al., 2022; Wei et al., 2022, 2023), connecting the factor model with the generative model, the variational autoencoder (VAE). These VAE-based models describe high-dimensional data (prices) to low-dimensional representations (factors), and learn factors self-adaptively. Although latent factor models have demonstrated substantial success in financial trading tasks, they still face several significant issues:
@@
  • CH2: Factor Inefficiency. In addition to the potential limitations of factors discussed in CH1, factors learned through VAEs exhibit three primary inefficiencies: i) Most latent factor models only focus on cross-sectional factors, neglecting the temporal perspective. ii) The continuous, multi-dimensional latent space may allow noise to overshadow meaningful factors. iii) The lack of independence among factors may lead to multicollinearity, with the homogeneity weakening the model’s adaptability to varying market conditions and limiting the effectiveness of factors in downstream investment tasks.
@@
In order to address the challenges, we propose a Spatio-Temporal factOR Model based on dual vector quantized variational autoencoders (VQ-VAE), named STORM, with the architecture shown in Figure 1. The model encodes factors as multi-dimensional embeddings, effectively capturing the complexity and nonlinearity inherent in financial data, and enhancing the capacity of factors to account for factor returns (CH1). Additionally, we develop a dual VQ-VAE architecture to capture cross-sectional and time-series features, considering both spatial and temporal perspectives [^2] We treat cross-sectional as spatial and time-series as temporal. These terms will be used interchangeably unless specified.. By integrating these features at both fine-grained and semantic levels, the model constructs more effective factors. Through diversity and orthogonality loss constraints, we ensure representation independence among learned representations (CH2). Furthermore, codebook embeddings act as cluster centers, serving as class tokens to categorize factor embeddings. This strategy provides clarity and transparency to the differentiation and selection process of factors (CH3). Specifically, our contributions are three-fold:
@@ 2.1. Factor Model
Factor model has attracted extensive research focus from the finance community. Financial experts have carefully selected a wide range of factors from macroeconomic (Sharpe, 1964), market changes (Fama and French, 1992), individual stock fundamentals (Greenblatt, 2010), and other perspectives to price assets in an attempt to capture excess returns from mispricing. Recently, latent factor models have emerged that improve investment returns by studying latent factors self-adaptively. Specifically, (Gu et al., 2021) learns latent factors depend on asset characteristics, (Duan et al., 2022) extracts cross-sectional factors by predicting future returns, and (Wei et al., 2023) learns market-augmented latent factors. However, these VAE-based latent factor models may suffer from a lack of robustness in representing low signal-to-noise ratio markets, leading to reduced reconstruction and generation capabilities. Furthermore, their effectiveness is constrained by an insufficient focus on the temporal perspective.
@@ 2.2. Financial Trading
Financial trading has attracted significant attention from both the finance and AI communities. Quantitative traders leverage mathematical models and algorithms to automatically identify trading opportunities (Sun et al., 2024). Within the FinTech domain, key research areas encompass portfolio management (PM), algorithmic trading (AT), order execution (Fang et al., 2021), market making (Spooner and Savani, 2020), etc. In this work, we focus on PM and AT, as they are fundamental to developing intelligent financial decision-making systems.
PM focuses on optimizing wealth allocation across assets (Li and Hoi, 2014). Its development has evolved from simple rule-based methods (Poterba and Summers, 1988) to prediction-based approaches using machine learning (Ke et al., 2017) and deep learning (Qin et al., 2017; Zhao et al., 2022) for predicting returns or price movements. Despite recent integration of reinforcement learning (RL) in PM (Zhang et al., 2024b), the latent factor models still predominantly rely on deep learning for market modeling and precise predictions.
AT, on the other hand, involves executing trades with algorithm-generated signals. Like PM, AT has seen a transition from rule-based and supervised learning techniques (Yang et al., 2020) to RL-driven approaches. In particular, RL-based methods (Sun et al., 2022) have gained prominence due to their ability to navigate complex sequential decision-making problems, further enhancing trade execution efficiency and adaptability in dynamic market environments.
@@ 2.3. Vector Quantized Variational Autoencoder
Vector Quantized Variational Autoencoder (VQ-VAE) has emerged as a fundamental framework in deep learning research, establishing methodological significance across core domains: representation learning, generative modeling, and time series analysis. Initially introduced by Van Den Oord et al., VQ-VAE has showcased its capabilities in processing high-dimensional data across diverse modalities, including images (Razavi et al., 2019), audio (Zhang et al., 2024a), and video (Yan et al., 2021), with a strong emphasis on learning discrete latent representations. Particularly noteworthy is the model’s demonstrated proficiency in capturing temporal dependencies, where recent studies like (Talukder et al., 2024) have validated its superior performance in modeling sequential patterns compared to conventional continuous latent space approaches, thereby advancing the development of robust generative architectures specifically optimized for temporal dynamics analysis.
@@ 3.1. Factor Model
The factor model (e.g., Arbitrage Pricing Theory (Ross, 1976)) is defined to study the relationship between expected asset returns and factor exposures, focusing on the cross-sectional differences in expected asset returns.
@@
According to (Duan et al., 2022), the factor model is formulated from a spatial perspective, primarily examining cross-sectional variations in expected returns across different assets, rather than capturing the temporal evolution of individual asset returns. Therefore, we expand it from the temporal perspective and introduce the general functional form of the latent factor model:
(2) $y_{i,t}=\alpha_{i,t}+\sum_{k=1}^{K}\bm{\beta}_{i,t}^{k}\mathbf{z}_{i,t}^{k}+\epsilon_{i,t}\text{,}\vspace{-0.1cm}$
@@ 3.2. Problem Formulation
To assess the effectiveness and generating ability of factors, we predict and evaluate the stock future returns. To show the adaptability of the latent factor model with multiple downstream tasks, we employ two financial trading tasks, i.e., portfolio management (Wang et al., 2021a; Zhang et al., 2024b; Ye et al., 2020) and algorithmic trading (Sun et al., 2022; Liu et al., 2020), to showcase the model’s investment capabilities.
@@
  • Portfolio management task aims to construct an optimal portfolio based on the predicted returns to test the profitability of the factor model. At each trading day $t$, we obtain the portfolio weight $\mathbf{w}_{t}=[w_{t}^{0},w_{t}^{1},...,w_{t}^{N-1}]\in\mathbb{R}^{N}$ for $N$ stocks. Each stock $i$ is assigned a weight $w_{t}^{i}$ representing its portfolio proportion within the portfolio. To ensure full capital deployment, the allocation must satisfy the budge constraint $\sum_{i=0}^{N-1}w_{t}^{i}=1$.
  • Algorithmic trading task aims to execute buy, hold, and sell actions based on predicted asset states to balance returns and risks. We formulate it as a Markov Decision Process (MDP) under reinforcement learning scenario following (Chen et al., 2023; Zhang et al., 2024c). The MDP is constructed by a 5-tuple $(\mathit{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma)$, where $\mathit{S}$ and $\mathcal{A}$ are sets of states and actions respectively, $\mathcal{T}:\mathit{S}\times\mathcal{A}\times\mathit{S}\rightarrow[0,1]$ is the state transition function, $\mathcal{R}:\mathit{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function where $\mathbb{R}$ is a continuous set of possible rewards, and $\gamma\in[0,1)$ is the discount factor. The goal is to learn a policy $\mathcal{\pi}:\mathit{S}\rightarrow\mathcal{A}$ that maximizes the expected discounted cumulative reward $\mathbb{E}[\sum^{\infty}_{t=0}\gamma^{t-1}r_{t}]$. At each trading day $t$ for specific stock $i$, the agent takes action $a_{i,t}:=\{a_{i,t}^{b},a_{i,t}^{h},a_{i,t}^{s}\}\in\mathcal{A}$ which means buy, hold and sell, according to the current environment $s_{t}\in\mathcal{S}$.
@@ 3.3. Vector Quantized Variational Autoencoder
VQ-VAE(Van Den Oord et al., 2017) is different from VAE (Kingma and Welling, 2013) as it quantizes the observed data into a discrete token sequence. The VQ-VAE contains three parts, encoder, decoder, and discrete codebook, denoted by $\phi_{\text{enc}}$ with parameter $\theta_{\text{enc}}$, $\phi_{\text{dec}}$ with parameter $\theta_{\text{dec}}$ and $\mathcal{C}=\{(k,e_{k}\in R^{H})\}_{k=1}^{K}$ with parameter $\{e_{k}\in R^{H}\}_{k=1}^{K}$ respectively. Among them, $K$ is the size of the codebook, $H$ is the dimension of the vector, and $e_{k}$ is the $k$-th vector of the latent embedding space. Given an input $x$, the encoder $\phi_{\text{enc}}$ firstly encodes it into a set of continuous feature vectors $Z=\phi_{\text{enc}}(x)$ and quantization operation $q(\cdot)$ passes each feature vector $z\in Z$ through the discretization bottleneck following by selecting the nearest neighbor embedding in the codebook $\mathcal{C}$ as its discrete code sequence $Q_{z}$.
(3) $Q_{z}=q(z,\mathcal{C})=c,\text{where }c=\arg\min_{k\in[0,K-1]}||z-e_{k}||_{2}^{2}\text{.}\vspace{-0.1cm}$
@@ 4.1.1. Patching
In stock data analysis, it is essential to extract features from both temporal and spatial dimensions. Time-series features capture the dynamics of a stock over time, while cross-sectional features reveal the correlations among different stocks. To achieve this, we partition data into patches (Nie et al., 2022) within the TS and CS modules. In the TS module, the observed data $\mathbf{x}$ is divided along the stock number dimension, with each patch containing data for one stock over $p$ days. Similarly, in the CS module, the observed data $\mathbf{x}$ is divided along the time axis, with each patch containing features of all stocks for a single trading day. Combining information from multi-scale features allows the model to capture both temporal and spatial dependencies, enhancing the depth and accuracy of market insights.
@@ 4.1.2. TS and CS Encoders
Considering the Transformer’s (Vaswani et al., 2017) remarkable ability to capture both local and global dependencies via self-attention and patching strategies, it is particularly well-suited for long-term time-series forecasting (Nie et al., 2022). Therefore, we employ stacked transformer blocks as the encoders in both the TS and CS modules to capture complex time-series and cross-sectional patterns. This architecture enhances the accuracy and robustness of feature extraction, ultimately improving forecasting performance.
While transformers are widely applicable across domains such as natural language processing (NLP) and computer vision (CV), STORM’s flexible encoder-decoder framework allows for the integration of alternative architectures, such as long short-term memory (LSTM) (Wang et al., 2021b) and temporal convolutional networks (TCN) (Yu and Koltun, 2015).

### v4

@@ Abstract.
In financial trading, factor models are widely used to price assets and capture excess returns from mispricing. Recently, we have witnessed the rise of variational autoencoder-based latent factor models, which learn latent factors self-adaptively. While these models focus on modeling overall market conditions, they often fail to effectively capture the temporal patterns of individual stocks. Additionally, representing multiple factors as single values simplifies the model but limits its ability to capture complex relationships and dependencies. As a result, the learned factors are of low quality and lack diversity, reducing their effectiveness and robustness across different trading periods. To address these issues, we propose a Spatio-Temporal factOR Model based on dual vector quantized variational autoencoders, named STORM, which extracts features of stocks from temporal and spatial perspectives, then fuses and aligns these features at the fine-grained and semantic level, and represents the factors as multi-dimensional embeddings. The discrete codebooks cluster similar factor embeddings, ensuring orthogonality and diversity, which helps distinguish between different factors and enables factor selection in financial trading. To show the performance of the proposed factor model, we apply it to two downstream experiments: portfolio management on two stock datasets and individual trading tasks on six specific stocks. The extensive experiments demonstrate STORM’s flexibility in adapting to downstream tasks and superior performance over baseline models. Code is available in PyTorch [^1] https://github.com/DVampire/Storm.
@@
journalyear: 2026copyright: ccconference: Proceedings of the Nineteenth ACM International Conference on Web Search and Data Mining; February 22–26, 2026; Boise, ID, USAbooktitle: Proceedings of the Nineteenth ACM International Conference on Web Search and Data Mining (WSDM ’26), February 22–26, 2026, Boise, ID, USAdoi: 10.1145/3773966.3777972isbn: 979-8-4007-2292-9/2026/02ccs: Information systems Spatial-temporal systemsccs: Applied computing Electronic commerce
@@ 1. Introduction
In financial trading, factor models are fundamental tools for asset pricing and are widely used to predict asset returns. These models enhance pricing accuracy and risk management by identifying a set of key factors that explain excess returns. The well-known Fama-French three-factor model (fama1992cross) extends the Capital Asset Pricing Model (sharpe1964capital) with additional factors to capture market risk premiums. As financial theories evolve, researchers continue to discover new factors, broadening our understanding of market dynamics.
@@
Recently, we have witnessed the rise of latent factor models (gu2021autoencoder; duan2022factorvae; wei2022factor; wei2023hirevae), connecting the factor model with the generative model, the variational autoencoder (VAE). These VAE-based models describe high-dimensional data (prices) to low-dimensional representations (factors), and learn factors self-adaptively. Although latent factor models have demonstrated substantial success in financial trading tasks, they still face several significant issues:
@@
  • CH2: Factor Inefficiency. VAE-learned factors suffer from three inefficiencies: i) focusing mainly on cross-sectional factors while neglecting temporal information, ii) allowing noise in continuous latent spaces to overshadow meaningful signals, and iii) lacking independence among factors, which leads to multicollinearity and weak adaptability to varying market conditions.
@@
In order to address the challenges, we propose a Spatio-Temporal factOR Model based on dual vector quantized variational autoencoders (VQ-VAE), named STORM, with the architecture shown in Figure 1. Unlike traditional scalar-valued financial factors with clear economic interpretations, STORM learns high-dimensional latent vectors that, while less interpretable capture complexity and nonlinearity inherent in financial martket (CH1). Additionally, we develop a dual VQ-VAE architecture to capture cross-sectional and time-series features, considering both spatial and temporal perspectives [^2] We treat cross-sectional as spatial and time-series as temporal. These terms will be used interchangeably unless specified.. By integrating these features at both fine-grained and semantic levels, the model constructs more effective factors. Through diversity and orthogonality loss constraints, we ensure representation independence among learned representations (CH2). Furthermore, codebook embeddings act as cluster centers, serving as class tokens to categorize factor embeddings. This strategy provides clarity and transparency to the differentiation and selection process of factors (CH3). Specifically, our contributions are four-fold:
@@ 2.1. Factor Model
Factor model has attracted extensive research focus from the finance community. Financial experts have carefully selected a wide range of factors from macroeconomic (sharpe1964capital), market changes (fama1992cross), individual stock fundamentals (greenblatt2010little), and other perspectives to price assets in an attempt to capture excess returns from mispricing. Recently, latent factor models have emerged that improve investment returns by studying latent factors self-adaptively. Specifically, (gu2021autoencoder) learns latent factors depend on asset characteristics, (duan2022factorvae) extracts cross-sectional factors by predicting future returns, and (wei2023hirevae) learns market-augmented latent factors. However, these VAE-based latent factor models may suffer from a lack of robustness in representing low signal-to-noise ratio markets, leading to reduced reconstruction and generation capabilities. Furthermore, their effectiveness is constrained by an insufficient focus on the temporal perspective.
@@ 2.2. Financial Trading
Financial trading has attracted significant attention from both the finance and AI communities. Quantitative traders leverage mathematical models and algorithms to automatically identify trading opportunities (sun2024trademaster). Within the FinTech domain, key research areas encompass portfolio management (PM), algorithmic trading (AT), order execution (fang2021universal), market making (spooner2020robust), etc. In this work, we focus on PM and AT, as they are fundamental to developing intelligent financial decision-making systems.
PM focuses on optimizing wealth allocation across assets (li2014online). Its development has evolved from simple rule-based methods (poterba1988mean) to prediction-based approaches using machine learning (ke2017lightgbm) and deep learning (qin2017dual; zhao2022stock) for predicting returns or price movements. Despite recent integration of reinforcement learning (RL) in PM (zhang2024reinforcement), the latent factor models still predominantly rely on deep learning for market modeling and precise predictions.
AT, on the other hand, involves executing trades with algorithm-generated signals. Like PM, AT has seen a transition from rule-based and supervised learning techniques (yang2020qlib) to RL-driven approaches. In particular, RL-based methods (sun2022deepscalper) have gained prominence due to their ability to navigate complex sequential decision-making problems, further enhancing trade execution efficiency and adaptability in dynamic market environments.
@@ 2.3. Vector Quantized Variational Autoencoder
Vector Quantized Variational Autoencoder (VQ-VAE) has emerged as a fundamental framework in deep learning research, establishing methodological significance across core domains: representation learning, generative modeling, and time series analysis. Initially introduced by van2017neural(van2017neural), VQ-VAE has showcased its capabilities in processing high-dimensional data across diverse modalities, including images (razavi2019generating), audio (zhang2024codebook), and video (yan2021videogpt), with a strong emphasis on learning discrete latent representations. Particularly noteworthy is the model’s demonstrated proficiency in capturing temporal dependencies, where recent studies like (talukder2024totem) have validated its superior performance in modeling sequential patterns compared to conventional continuous latent space approaches, thereby advancing the development of robust generative architectures specifically optimized for temporal dynamics analysis.
@@ 3.1. Factor Model
The factor model (e.g., Arbitrage Pricing Theory (ross1976arbitrage)) is defined to study the relationship between expected asset returns and factor exposures, focusing on the cross-sectional differences in expected asset returns.
@@
According to (duan2022factorvae), the factor model is formulated from a spatial perspective, primarily examining cross-sectional variations in expected returns across different assets, rather than capturing the temporal evolution of individual asset returns. Therefore, we expand it from the temporal perspective and introduce the general functional form of the latent factor model:
(2) $y_{i,t}=\alpha_{i,t}+\sum_{k=1}^{K}\bm{\beta}_{i,t}^{k}\mathbf{z}_{i,t}^{k}+\epsilon_{i,t}\text{,}$
@@ 3.2. Problem Formulation
To assess the effectiveness and generating ability of factors, we predict and evaluate the stock future returns. To show the adaptability of the latent factor model with multiple downstream tasks, we employ two financial trading tasks, i.e., portfolio management (wang2021deeptrader; zhang2024reinforcement; ye2020reinforcement) and algorithmic trading (sun2022deepscalper; liu2020adaptive), to showcase the model’s investment capabilities.
@@
  • Portfolio management task aims to construct an optimal portfolio based on the predicted returns to test the profitability of the factor model. At each trading day $t$, the factor model outputs return predictions $\hat{\mathbf{y}}_{i,t+1}$ for each stock $i$, which are then used by a portfolio allocation strategy, e.g., TopK-Drop, to generate the portfolio weight vector $\mathbf{w}_{t}=[w_{t}^{0},w_{t}^{1},\dots,w_{t}^{N-1}]\in\mathbb{R}^{N}$, where $w_{t}^{i}$ represents the proportion of capital invested in stock $i$. To ensure full capital deployment, the allocation must satisfy the budget constraint $\sum_{i=0}^{N-1}w_{t}^{i}=1$.
  • Algorithmic trading task aims to execute buy, hold, and sell actions based on predicted asset states to balance returns and risks. We formulate it as a Markov Decision Process (MDP) under reinforcement learning scenario following (chen2023mtrader; zhang2024finagent). The MDP is constructed by a 5-tuple $(\mathit{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma)$, where $\mathit{S}$ and $\mathcal{A}$ are sets of states and actions respectively, $\mathcal{T}:\mathit{S}\times\mathcal{A}\times\mathit{S}\rightarrow[0,1]$ is the state transition function, $\mathcal{R}:\mathit{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function where $\mathbb{R}$ is a continuous set of possible rewards, and $\gamma\in[0,1)$ is the discount factor. The latent factor embeddings $\mathbf{Z}$ generated by the factor model are integrated into the observation set that constitutes the state space $\mathcal{S}$, enabling the agent to incorporate predictive market signals. The goal is to learn a policy $\mathcal{\pi}:\mathit{S}\rightarrow\mathcal{A}$ that maximizes the expected discounted cumulative reward $\mathbb{E}[\sum^{\infty}_{t=0}\gamma^{t-1}r_{t}]$. At each trading day $t$ for specific stock $i$, the agent takes action $a_{i,t}:=\{a_{i,t}^{b},a_{i,t}^{h},a_{i,t}^{s}\}\in\mathcal{A}$ which means buy, hold and sell, according to the current environment $s_{t}\in\mathcal{S}$.
@@ 3.3. Vector Quantized Variational Autoencoder
VQ-VAE(van2017neural) is different from VAE (kingma2013auto) as it quantizes the observed data into a discrete token sequence. The VQ-VAE contains three parts, encoder, decoder, and discrete codebook, denoted by $\phi_{\text{enc}}$ with parameter $\theta_{\text{enc}}$, $\phi_{\text{dec}}$ with parameter $\theta_{\text{dec}}$ and $\mathcal{C}=\{(k,e_{k}\in R^{H})\}_{k=1}^{K}$ with parameter $\{e_{k}\in R^{H}\}_{k=1}^{K}$ respectively. Among them, $K$ is the size of the codebook, $H$ is the dimension of the vector, and $e_{k}$ is the $k$-th vector of the latent embedding space. Given an input $x$, the encoder $\phi_{\text{enc}}$ firstly encodes it into a set of continuous feature vectors $Z=\phi_{\text{enc}}(x)$ and quantization operation $q(\cdot)$ passes each feature vector $z\in Z$ through the discretization bottleneck following by selecting the nearest neighbor embedding in the codebook $\mathcal{C}$ as its discrete code sequence $Q_{z}$.
[…]
@@ 4.1.1. Patching
In stock data analysis, it is essential to extract features from both temporal and spatial dimensions. Time-series features capture the dynamics of a stock over time, while cross-sectional features reveal the correlations among different stocks. To achieve this, we partition data into patches (nie2022time) within the TS and CS modules. In the TS module, the observed data $\mathbf{x}$ is divided along the stock number dimension, with each patch containing data for one stock over $p$ days (as illustrated in Figure 2). Similarly, in the CS module, the observed data $\mathbf{x}$ is divided along the time axis, with each patch containing features of all stocks for a single trading day. Combining information from multi-scale features allows the model to capture both temporal and spatial dependencies, enhancing the depth and accuracy of market insights.
@@ 4.1.2. TS and CS Encoders
Considering the Transformer’s (vaswani2017attention) remarkable ability to capture both local and global dependencies via self-attention and patching strategies, it is particularly well-suited for long-term time-series forecasting (nie2022time). Therefore, we employ stacked transformer blocks as the encoders in both the TS and CS modules to capture complex time-series and cross-sectional patterns. This architecture enhances the accuracy and robustness of feature extraction, ultimately improving forecasting performance.
While transformers are widely applicable across domains such as natural language processing (NLP) and computer vision (CV), STORM’s flexible encoder-decoder framework allows for the integration of alternative architectures, such as long short-term memory (LSTM) (wang2021clvsa) and temporal convolutional networks (TCN) (yu2015multi).
An encoder–decoder configuration comprising 4 and 2 blocks, respectively, each with 4 and 8 attention heads, remains relatively lightweight, avoiding excessive computational cost and a substantial increase in model complexity.